λ_{obs} = λ_{source} \begin{matrix}√ & \overline{1 + β} \\ √ & \overline{1 − β} \end{matrix}  =  λ_{source} \begin{matrix}√ & \overline{1 + (v/c)} \\ √ & \overline{1 − (v/c)} \end{matrix}

λ_{obs} ≃ λ_{source} (1 + β)  =  λ_{source} (1 + (v/c))

velocity (km/s) = H distance (Mpc)

z = \frac{λ_{obs}}{λ_{source}} − 1

β ≡ \frac{v}{c}